What is the height of a trapezoid with an area of 225 square centimeters and bases of 22 centimeters and 28 centimeters
Answer by Natolino1983(23)

Area of Trapezoid = average between (bases) × height
Average is (22+28)/2 = 25 cm.
And Area = 225 cm^2
So height = Area/Average = 225/25 = 9 cm